40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 13 21:30:55 2020, shruti, Update, ALS, ALS OOL noise with arms locked 20200213_ALS.pdf
    Reply  Fri Feb 14 14:02:13 2020, shruti, Update, ALS, ALS OOL noise with arms locked 819472_10.pdf225642_10.pdf693500_10.pdf
       Reply  Fri Feb 14 14:52:41 2020, gautam, Update, ALS, ALS OOL noise with arms locked 
          Reply  Tue Feb 18 18:14:59 2020, shruti, Update, ALS, ALS OOL noise with arms locked ALSNoise.pdf
             Reply  Wed Feb 19 22:20:22 2020, rana, Update, ALS, ALS OOL noise with arms locked 
             Reply  Fri Feb 21 10:59:08 2020, shruti, Update, ALS, PDH error signals? IMG_0082.jpgIMG_0083.jpgIMG_0084.jpgIMG_0085.jpg
                Reply  Fri Feb 21 13:02:53 2020, Koji, Update, ALS, PDH error signals? 
                   Reply  Fri Feb 21 20:44:18 2020, shruti, Update, ALS, ALS OOL noise and PDH ALSNoise.pdfIMG_0086.jpg
                      Reply  Sun Feb 23 18:15:22 2020, rana, Update, ALS, ALS OOL noise and PDH 
Message ID: 15211     Entry time: Thu Feb 13 21:30:55 2020     Reply to this: 15213
Author: shruti 
Type: Update 
Category: ALS 
Subject: ALS OOL noise with arms locked 

[Meenakshi, Gautam, Shruti]

Summary:

- We initially aligned the arm cavities to get the green lasers locked to them. For the X arm cavity, we tweaked the ITMX and ETMX pitch and yaw and toggled the X green shutter until we saw something like a TEM00 mode on the monitor and a reasonable transmitted power.

- With the LSC servo enabled, the IR light also became resonant with the cavities.

- Then we measured the noise in different configurations. Attachment 1 shows the the ALS OOL (in the IR beat signal) noise with the arms locked inidividually via PDH.


The script for plotting the ALS beat frequency noise is:

users/Templates/ALS/ALS_outOfLoop_Ref.xml
Attachment 1: 20200213_ALS.pdf  75 kB  | Hide | Hide all
20200213_ALS.pdf
ELOG V3.1.3-